# LLaVA-UHD v2: an MLLM Integrating High-Resolution Feature Pyramid via Hierarchical Window Transformer

### v1

@@ 3.3 Hierarchical Window Attention
The hierarchical nature of feature pyramids necessitates an effective approach for compressing features at varying resolutions while maintaining cross-level spatial alignment.
Hierarchical window generation. Inspired by object detection [86, 114], we employ the RoI-align [36] to sample key features while preserving the spatial locality across feature maps at different levels, as illustrated in Fig. 4. Specifically, we begin by uniformly dividing the feature maps of each level into $N\times N$ windows, whose widths and heights are float-point values $(\frac{W}{N},\frac{H}{N})$ rather than integers. Especially, windows that share the same “anchor” points across levels form a set of hierarchical bounding boxes, represented by the top-left and bottom-right coordinates $\{\mathcal{R}^{l}_{i,j}\in\mathbb{R}^{1\times 4},i,j\in 0,1,2...N-1,l=0,1,2\}$, where $l$ indicates the feature level and $(i,j)$ the 2D indices. To mitigate the size distortion of feature maps caused by the discrepancy between the aspect ratio of RoI-aligned feature maps and the original feature maps, we define a pooling score to evaluate this discrepancy:
@@
where $(r_{w},r_{h})$ denotes the width and height of pooled features. By maximizing the score $S$, we select the optimal grid size $(r^{*}_{w},r^{*}_{h})$ from a set of predefined proposals $\{(3,3),(2,3),(3,2),(2,4),(4,2)\}$. Subsequently, we perform RoI-align with the generated windows to sample the key feature maps for the following attention operation.
Cross-window feature querying. To compress the multi-level feature pyramid $\{\mathcal{F}^{l},l=0,1,2\}$ of a given image slice $\mathcal{I}$, we initialize a set of queries $\{\mathcal{Q}_{i,j}\in\mathbb{R}^{1\times C},i,j\in 0,1,2...N-1\}$, each of which corresponds to a set of hierarchical windows $\{\mathcal{R}^{l}_{i,j},l=0,1,2\}$, in Fig. 4. Given each query vector $\mathcal{Q}_{i,j}$, we prepare the key vector $\mathcal{K}^{l}_{i,j}\in\mathbb{R}^{(r^{*}_{x}\cdot r^{*}_{y})\times C}$ of $l$-th level as
@@
where $\phi^{l}$ is a level-wise positional embedding. The key vectors $\mathcal{K}^{l}_{i,j}$ are then concatenated along the length axis to form the final key vector $\mathcal{K}_{i,j}\in\mathbb{R}^{(3\cdot r^{*}_{x}\cdot r^{*}_{y})\times C}$ for each query $\mathcal{Q}_{i,j}$. The corresponding value vector $\mathcal{V}_{i,j}$ is obtained in the same manner, but without the level-wise positional embedding. Thus, the cross-attention can be performed as
$\mathcal{Q^{*}}_{i,j}=\text{CrossAttn}(\mathcal{Q}_{i,j}+\varphi_{i,j},\mathcal{K}_{i,j}+\phi_{i,j},\mathcal{V}_{i,j}),$ (7)
where $\mathcal{Q^{*}}_{i,j}$ denotes the updated query, and $\varphi,\phi$ is the 2D spatial position embedding of query and key vector, respectively. In the end, we concatenate all the updated query $\mathcal{Q^{*}}_{i,j}$ into a feature map $\mathcal{P}\in\mathbb{R}^{N\times N\times C}$ to represent the visual token of image $\mathcal{I}$.
@@ 3.4 Spatially-consistent Token Organization
Due to the varying partitions across different images, effectively organizing and conveying the structure of slices to the LLM is crucial for a more accurate understanding of the image. Prior studies [31, 60, 103] employ spatial tokens (e.g., “\$\backslash n$” or “,”) to denote the relative positioning of image slices. While such approaches provide cues regarding the global arrangement of each slice, they overlook the intrinsic spatial relationships among individual visual tokens. For instance, two horizontally adjacent tokens in the image, located in different slices, may become significantly separated in a 1D arrangement. To address this, we leverage our Hiwin transformer, which preserves 2D spatial consistency with the original image, and amalgamate these 2D feature maps into a large 2D feature map according to the slicing configuration, in Fig. 2. A newline token “\$\backslash n$” is then inserted between rows, followed by flattening them into the 1D arrangement. Additionally, a comma “,” is placed between the large feature maps and the features of the overview image to distinguish between different views of the image.
@@ 4 Experiment
In this section, we present an empirical evaluation of LLaVA-UHD v2. We begin with a comprehensive outline of the implementation details of our model, followed by a comparative analysis of its performance across widely recognized benchmarks against competitive counterparts. Finally, we provide an in-depth ablation to further elucidate the capabilities and behavior of LLaVA-UHD v2.
@@ 4.1 Implementation Details
Model Setting. We adopt LLaVA-UHD [31] as the baseline method. Specifically, we employ CLIP-ViT-L/14-336 as the visual encoder, Vicuna-7B [21] as the language model, and our proposed Hiwin Transformer as the vision-language projector. We set the maximum slice number to 6 to cover a range of aspect ratios and image resolutions. The number of learnable local queries is set as 144 ($i.e.,$$N=12$). LLaVA-UHD v2 consists of a three-stage training process, as outlined below.
Stage 0: JBU module pre-training. Before employing the JBU module in MLLM, we pre-train it on the MS-COCO [58] dataset with a global batch of 16 on 8$\times$×A100. We leverage Adam optimizer with a learning rate of 1e-3 for 2000 steps. During this stage, the parameters of CLIP-ViT are frozen.
Stage 1: MLLM pre-training. In this stage, the parameters of the visual encoder, JBU module, and LLM are frozen. We only fine-tune the parameters within the hierarchical window attention of the HiWin Transformer using LLaVA-Pretrain [63] dataset for 1 epoch. We employ the AdamW optimizer with a learning rate of 1e-3 and a cosine learning rate schedule. The global batch size is set to 256. Note that during this stage, we only encode the overview image, excluding the slices, for efficiency.
@@ 4.2 Experimental Setting
Benchmarks. Extensive benchmarks are used to analyze the effectiveness of our modules. We categorize these benchmarks into the following groups: (1) General VQA benchmarks including MME [25], MMB [65], SEED-Image [50] and GQA [38]; (2) Knowledge-based VQA benchmarks including MMMU-val [108], Science-QA [72], AI2D [44]; (3) OCR-based VQA benchmarks including ChartQA [77], OCR-Bench [66], TextVQA [94] and DocVQA [78]; (4) Visual spatial understanding benchmarks such as RealWorldQA [1] and RefCOCOs [107] (used in ablation studies); (5) High-resolution image perception benchmarks like HR-Bench(4K) [100].
Evaluation Protocols. Beyond benchmark evaluations, we report additional metrics for comprehensive analysis, including: (1) overall volume of training data, (2) maximum supported image resolution for each method, and (3) computational cost of the entire MLLM at maximum resolution.
Counterparts. We compare our model with the advanced MLLM counterparts. (1) General MLLMs like Qwen-VL [10], MiniGPT-v2 [16], Honey-bee [15], mPLUG-Owl2 [105], VILA [56] and LLaVA-1.5 [63]. (2) High-resolution MLLMs including UReader [104], Monkey [55], LLaVA-HR [73] ,LLaVA-Next [62]. (3) Mixture of visual experts such as  [73], SPHINX-series [60, 28] and Mini-Gemini [7].
Table 1: Main performance on popular benchmarks. For a fair comparison, we only report the method using 7B level LLM ($e.g.$, Vicuna-7B). #Data denotes the volume of overall data during MLLM pre-training and supervised fine-tuning. “MaxRes.” is the maximum accessible resolution of MLLM. “Avg.”: average results of 13 benchmarks. “VQAD: DocVQA. “BenchOCR”: OCR-Bench. “VQAC”: ChartQA. “VQAT”: TextVQA. “SQA”: Science-QA. “MMMUv”: MMMU-val. “SEEDI”: SEED-Image. “MMEP”: perception sub-set of MME. “RWQA”: RealWorldQA. “BenchHR”: HR-Bench.
@@
Method #Data MaxRes. #FLOPs. Avg. VQAD BenchOCR VQAC VQAT AI2D SQA MMMUv GQA SEEDI MMB MMEP RWQA BenchHR
Qwen-VL [10] 1.45B 448$\times$448 4.0T 56.9 62.6 48.8 66.3 61.5 57.7 68.2 35.9 57.5 65.4 61.8 74.4 49.3 30.5
MiniGPT-v2 [16] 326M 448$\times$448 4.0T - - 15.7 - - - - - 60.3 - - - - -
mPLUG-Owl2 [105] 401M 448$\times$448 1.7T - - - - 58.2 - 68.7 - 56.1 57.8 64.5 72.5 - -
UReader [104] 86M 896$\times$1120 20.3T - 65.4 - 59.3 57.6 - - - - - - - - -
LLaVA-1.5 [63] 1.22M 336$\times$336 8.0T 49.0 21.8 31.8 17.8 45.5 55.5 66.8 37.0 62.0 65.8 66.5 75.3 54.8 36.1
SPHINX-2k [60] 1.01B 762$\times$762 42.2T - - - - 61.2 - 70.6 - 63.1 71.6 65.9 73.6 - -
SPHINX-X [28] 15.3M 448$\times$448 21.3T - 56.3 - 39.7 58.1 63.0 70.4 - 56.2 68.8 57.9 63.0 - -
LLaVA-HR [73] 1.22M 1024$\times$1024 24.3T - - - - 67.1 - 65.1 - 64.2 64.2 - 77.7 - -
VILA [56] 51M 336$\times$336 8.2T - - - - 64.4 - 68.2 - 62.3 61.1 68.9 76.7 - -
Honey-bee [15] 52.5M 336$\times$336 2.6T - - - - - - - 35.3 - 64.5 70.1 79.2 - -
Mini-Gemini [54] 3.0M 672$\times$672 54.6T 59.4 61.9 47.7 47.4 65.2 68.2 69.6 36.8 64.5 66.9 65.8 77.3 51.1 50.1
Monkey [55] 1.40B 896$\times$1344 28.0T 59.2 66.5 51.4 65.1 67.6 62.6 69.4 38.9 60.7 64.3 59.8 73.6 51.6 38.0
LLaVA-Next [62] 1.34M 672$\times$672 44.4T 61.0 63.6 53.2 54.3 64.9 67.0 70.1 35.8 64.2 70.2 67.4 76.0 57.8 47.9
LLaVA-UHD v2 (ours) 1.42M 1008$\times$672 17.5T 63.2 68.1 53.9 64.5 67.6 70.5 71.3 38.2 65.4 70.0 68.2 74.7 58.2 51.5
Table 2: Ablation studies of modules in our proposed method. “HFP” is the abbreviation of high-resolution feature pyramid. “$\Delta$” denotes the overall improvement compared to the baseline. REC reports the average accuracy of RefCOCO/g/+.
@@ 4.3 Main Performance
Table 1 showcases a comparative analysis of our proposed LLaVA-UHD v2 against state-of-the-art MLLMs across 13 widely recognized benchmarks. To facilitate a comprehensive evaluation, we have selected 5 prominent open-source models, and assessed them on all 13 benchmarks, to serve as our reference points. (1) LLaVA-UHD v2 outperforms current counterparts. Our model exhibits a significant enhancement in average performance, achieving an improvement of 2.2% compared to LLaVA-Next, among other models. Specifically, compared with general models (such as Qwen-VL and LLaVA-1.5) and high-resolution MLLMs (like Monkey and LLaVA-Next), LLaVA-UHD v2 demonstrates consistent improvements across various tasks, including general VQA (e.g., 65.4% on GQA), ultra-high-resolution image perception (e.g., 51.5% on HR-Bench). Notably, LLaVA-UHD v2 surpasses OCR-centric models like UReader on DocVQA (68.1% vs. 65.4%) and outperforms those with multiple experts (such as SPHINX-2k and Mini-Gemini), achieving superior performance on general tasks like MMB (68.2%) and SEED (70.0%). These results highlight the value of rich semantics derived from multi-level visual granularity in enhancing both the understanding and perceptual capabilities of MLLMs. (2) LLaVA-UHD v2 indicates efficiency on data utilization and computation. Compared to LLaVA-Next and Mini-Gemini, both operating at the 672$\times$672 resolution, LLaVA-UHD v2 supports 1.5 times the resolution ($i.e$., 672$\times$1008) and achieves superior performance with less than 40% of the computational cost. Furthermore, in contrast to Honey-bee and VILA, which utilize 52.5M and 51M data samples respectively, LLaVA-UHD v2 attains comparable or superior performance using only $\sim$2.8of the data, demonstrating the data efficiency of our model, making it highly suitable for low-cost exploratory research in the academic community.
Figure 5: Performance on different visual tasks with JBU module and vanilla bilinear interpolation. “OCR” denotes the optical character recognition, “Seg” the Linear probing semantic segmentation, and “Cls” the fine-grained classification on SUB-200.
Table 3: Comparison of different methods for feature pyramid construction. “ConvNext” means we replace the CILP-ViT with CLIP-ConvNext [68] as visual encoder and directly use the feature maps from multiple stages as the final hierarchical feature pyramid.

### v2

@@ 3.3 Hierarchical Window Attention
The hierarchical nature of ISP necessitates an effective approach for compressing features at varying resolutions while maintaining cross-level spatial alignment.
Hierarchical window generation. Inspired by object detection [92, 122], we utilize the RoI-align [40] to sample key features to keep the spatial locality of cross-level feature maps, in Fig. 4. Specifically, we first uniformly divide feature maps of each level into $N\times N$ windows, whose widths and heights are float-point values $(\frac{W}{N},\frac{H}{N})$ rather than integers. Windows share the same “anchor" point form a set of hierarchical bounding boxes (coordinates of top-left and bottom-right) $\{\mathcal{R}^{l}_{i,j}\in\mathbb{R}^{1\times 4},i,j\in 0,1,2...N-1,l=0,1,2\}$, where $l$ is the feature level and $(i,j)$ the 2D index. To mitigate the size distortion of feature maps caused by a difference between the aspect ratio of RoI-aligned feature maps and the image, we define a pooling score to evaluate this difference:
@@
where $(r_{w},r_{h})$ denotes the width and height of pooled features. By maximizing the score $S$, we select the optimal grid size $(r^{*}_{w},r^{*}_{h})$ from pre-defined proposals $\{(3,3),(2,3),(3,2),(2,4),(4,2)\}$. Then, we carry out the RoI-align with the generated windows to sample the key feature maps for the following attention operation.
Cross-scale window querying. To compress the ISP $\{\mathcal{F}^{l},l=0,1,2\}$ of one image or slice $\mathcal{I}$, we initialize a set of queries $\{\mathcal{Q}_{i,j}\in\mathbb{R}^{1\times C},i,j\in 0,1,2...N-1\}$, each of which corresponds to a set of hierarchical windows $\{\mathcal{R}^{l}_{i,j},l=0,1,2\}$, in Fig. 4. Regarding each query vector $\mathcal{Q}_{i,j}$, we prepare the key vector $\mathcal{K}^{l}_{i,j}\in\mathbb{R}^{(r^{*}_{w}\cdot r^{*}_{h})\times C}$ of $l$-th level as
@@
where $\phi^{l}$ is a level positional embedding. We then concatenate the $\mathcal{K}^{l}_{i,j}$ in length axis and form the final key vector $\mathcal{K}_{i,j}\in\mathbb{R}^{(3\cdot r^{*}_{w}\cdot r^{*}_{h})\times C}$ for each $\mathcal{Q}_{i,j}$. The corresponding value vector $\mathcal{V}_{i,j}$ is obtained in the same way yet without the level positional embedding. Thus, the cross-attention can be performed as
$\mathcal{Q^{*}}_{i,j}=\text{CrossAttn}(\mathcal{Q}_{i,j}+\varphi_{i,j},\mathcal{K}_{i,j}+\zeta_{i,j},\mathcal{V}_{i,j}),$ (5)
where $\mathcal{Q^{*}}_{i,j}$ denotes the updated query, and $\varphi,\zeta$ is the 2D spatial position embedding of query and key vector, respectively. In the end, we concatenate all the $\mathcal{Q^{*}}_{i,j}$ into a feature map $\mathcal{P}\in\mathbb{R}^{N\times N\times C}$ to represent the visual token of $\mathcal{I}$.
Figure 4: The flowchart of hierarchical window attention. We initialize a set of learnable queries to attend to local regions. Feature maps from the ISP are processed by a set of cross-scale windows, forming hierarchical and local-aware features at different levels. The features are then concatenated along the length axis, to serve as the key and value for the learnable queries. The output is condensed visual tokens rich in diverse and local-aware semantics.
@@ 3.4 Spatially-consistent Token Organization
Due to the varying partitions of different images, organizing and conveying the structure of slices to the LLM facilitates a more accurate understanding of the image. Prior studies [35, 64, 110] utilize spatial tokens (e.g., “\$\backslash n$” or “,”) to denote the relative positioning of image slices. While such approaches provide cues regarding the global arrangement of each slice, they overlook the intrinsic spatial relationships among individual visual tokens. For instance, two horizontally adjacent tokens in the image, located in different slices, may become significantly separated in a 1D arrangement. Leveraging our Hiwin transformer, which preserves the 2D spatial consistency with the original image, we amalgamate these 2D feature maps into a large 2D feature map according to the slicing configuration, in Fig. 2.
@@ 4 Experiment
In this section, we conduct an empirical evaluation of LLaVA-UHD v2. We begin with a comprehensive outline of the implementation details of our model, followed by a comparative analysis of its performance across widely recognized benchmarks against competitive counterparts. Finally, we provide an in-depth ablation to further elucidate the capabilities and behaviors of LLaVA-UHD v2.
@@ 4.1 Implementation Details
Stage 1: MLLM pre-training. In this stage, the parameters of the visual encoder, pre-trained VDIM, and LLM are frozen. We only fine-tune the parameters within the hierarchical window attention of the HiWin transformer using LLaVA-Pretrain [67] for 1 epoch with a global batch size of 256. We employ the AdamW optimizer and a cosine learning rate scheduler. The learning rate is 1e-3 for Vicuna-7B, 2e-4 for Vicuna-13B, and Qwen2-7B. Note that, in this stage, we only encode the overview image without the slices for efficiency.
Stage 2: MLLM supervised fine-tuning. In this stage, we fine-tune all parameters except those in the VDIM. The learning rate is set to 2e55-5- 5 with a batch size of 128. To manage training costs, we use 825k data for analysis and ablation studies, including LLaVA-mix665k [67] and 160k from Ureader [111]. For comparison with advanced MLLMs, we balance our data distribution and introduce an 858k-mixed dataset, which was detailed in supplementary.
@@ 4.2 Experimental Setting
Benchmarks. Extensive benchmarks are used to analyze the effect of our modules. We categorize these benchmarks into the following folds: (1) General VQA benchmarks including MME [28], MMB [69], SEED-Image [54], GQA [42], MMStar [20] and HallusionBench: [34] ; (2) Knowledge-based VQA benchmarks including MMMU-val [115], Science-QA [78], AI2D [48], MathVista [79]; (3) OCR-based VQA benchmarks including ChartQA [84], OCR-Bench [70], TextVQA [100] and DocVQA [85]; (4) Visual spatial understanding benchmarks such as RealWorldQA [1] and RefCOCOs [114]; (5) High-resolution image perception benchmarks like HR-Bench(4K) [105].
Evaluation Protocols. Beyond benchmark evaluations, we report additional metrics for comprehensive analysis: (1) overall volume of training data, (2) maximum supported image resolution for each method, and (3) computation cost of the entire MLLM at maximum resolution.
Counterparts. We compare our model with the advanced MLLM counterparts. (1) General MLLMs like Honey-bee [16], Dense Connector [109], VILA [60] and LLaVA-1.5 [67]. (2) High-resolution MLLMs including Monkey [59], LLaVA-Next [66], PIIP-LLaVA [107], SliME-Llama3-8B [118], DeepseekVL-7B [76] and Token-Packer [57]. (3) Mixture of visual experts such as LLaVA-HR [80], SPHINX-series [64, 31] , MG-LLaVA [120] and Mini-Gemini [8]. (4) OCR-centric MLLMs including UReader [111] and TextMonkey [71].
Table 1: Main performance on popular benchmarks. #Data denotes the volume of overall data during MLLM pre-training and supervised fine-tuning. “MaxRes." is the maximum accessible resolution of MLLM. “VQAD": DocVQA. “BenchOCR": OCR-Bench. “VQAC": ChartQA. “VQAT": TextVQA. “SQA": Science-QA. “MMMUv": MMMU-val. “Math.": MathVista. “SEEDI": SEED-Image. “MMEP": perception sub-set of MME. “RWQA": RealWorldQA. “BenchHR": HR-Bench.
@@
Method LLM #Data MaxRes. #FLOPs. VQAD BenchOCR VQAC VQAT AI2D SQA MMMUv Math. GQA SEEDI MMB MMEP MMStar RWQA BenchHR
mPLUG-Owl2 [112] Llama2-7B 401M 448$\times$448 1.7T - - - 58.2 - 68.7 - 25.5 56.1 57.8 64.5 72.5 34.8 - -
UReader [111] Llama2-7B 86M 896$\times$1120 20.3T 65.4 - 59.3 57.6 - - - - - - - - - - -
VILA [60] Llama2-7B 51M 336$\times$336 8.2T - - - 64.4 - 68.2 - - 62.3 61.1 68.9 76.7 - - -
SPHINX-2k [64] Llama2-Accessory 1.01B 762$\times$762 42.2T - - - 61.2 - 70.6 - - 63.1 71.6 65.9 73.6 - - -
SPHINX-X [31] Llama2-Accessory 15.3M 448$\times$448 21.3T 56.3 - 39.7 58.1 63.0 70.4 - - 56.2 68.8 57.9 63.0 - - -
LLaVA-HR [80] Vicuna-7B 1.22M 1024$\times$1024 24.3T - - - 67.1 - 65.1 - - 64.2 64.2 - 77.7 - - -
Honey-bee [16] Vicuna-7B 52.5M 336$\times$336 2.6T - - - - - - 35.3 - - 64.5 70.1 77.2 - - -
Mini-Gemini [58] Vicuna-7B 3.0M 672$\times$672 54.6T 61.9 47.7 47.4 65.2 68.2 69.6 36.8 - 64.5 66.9 65.8 77.3 - 51.1 50.1
Monkey [59] Vicuna-7B 1.40B 896$\times$1344 28.0T 66.5 51.4 65.1 67.6 62.6 69.4 38.9 33.5 60.7 64.3 59.8 73.6 37 51.6 38.0
LLaVA-1.5 [67] Vicuna-7B 1.22M 336$\times$336 8.0T 21.8 31.8 17.8 45.5 55.5 66.8 37.0 25.5 62.0 65.8 66.5 75.3 33.1 54.8 36.1
LLaVA-Next [66] Vicuna-7B 1.34M 672$\times$672 44.4T 63.6 53.2 54.3 64.9 67.0 70.1 35.8 34.6 64.2 70.2 67.4 76.0 37.6 57.8 47.9
Token-Packer [57] Vicuna-7B 2.7M 1008$\times$1008 13.1T 60.2 45.2 - 68.0 - - 35.4 - - 67.4 74.5 - - -
TextMonkey [71] Vicuna-7B 1.45B 448$\times$448 4.0T 66.7 - 59.9 64.3 - - - - - - - - - - -
LLaVA-1.5 [67] Vicuna-13B 1.22M 336$\times$336 15.1T - - - 61.3 - 71.6 - - 63.3 61.6 67.7 76.5 - - -
LLaVA-Next [66] Vicuna-13B 1.34M 672$\times$672 67.0T - 53.7 61.4 67.1 - 73.6 36.2 35.3 65.4 71.9 70.0 76.5 40.4 57.6 -
LLaVA-1.5-Qwen [7] Qwen2-7B 1.22M 336$\times$336 8.2T - - - - 64.9 - 40.7 33.6 62.7 69.4 72.0 76.0 - - -
Dense Connector [109] Llama3-8B 1.22M 384$\times$384 11.6T - - - - - 75.2 40.4 28.6 65.1 - 74.4 - - - -
LLaVA-LLaMA3 [24] Llama3-8B 1.22M 336$\times$336 8.7T - - - - - 73.3 36.8 - 63.5 - 68.9 - - - -
PIIP-LLaVA [107] Llama3-8B 1.22M 1024$\times$1024 36.0T - - - 67.1 - 68.3 - - 63.9 69.4 67.0 - - - -
MG-LLaVA [120] Llama3-8B 2.5M 768$\times$768 33.0T - - - 67.3 - 70.8 - - - 69.4 72.1 - - - -
SliME [118] Llama3-8B 2.0M 2016$\times$2016 62.0T - - - 64.8 - - 41.2 - 63.9 - 75.0 - - - -
DeepseekVL-7B [76] DeepseekLLM-7B - 1024$\times$1024 - - 45.6 - - - - 36.6 36.1 - 70.4 73.2 - 37.1 - -
LLaVA-UHD v2 Vicuna-7B 1.42M 1008$\times$672 17.5T 68.1 53.9 64.5 67.6 70.5 71.3 38.2 34 65.4 70.0 68.2 74.7 40.2 58.2 51.5
LLaVA-UHD v2 Vicuna-13B 1.42M 1008$\times$672 26.4T 68.2 55.6 67.4 70.0 72.4 73.3 37.7 35.2 66.0 71.1 70.3 73.1 42.0 59.6 55.3
LLaVA-UHD v2 Qwen2-7B 1.42M 1008$\times$672 14.3T 72.9 57.7 70.4 70.6 75.5 76.9 43.3 39.1 65.1 73.6 77.1 78.8 49.4 64.6 59.9
Table 2: Ablation studies of modules in our proposed method. “$\Delta$" denotes the overall improvement compared to the baseline. REC reports the average accuracy of RefCOCO/g/+.
@@ 4.3 Main Performance
Table 1 showcases a comparative analysis of our proposed LLaVA-UHD v2 against state-of-the-art MLLMs across 15 widely recognized benchmarks. (1) LLaVA-UHD v2 outperforms current counterparts. Compared with general models (such as LLaVA-1.5, Dense Connector) and high-resolution MLLMs (like PIIP-LLaVA, SliME-Llama3-8 and DeepseekVL-7B), LLaVA-UHD v2 demonstrates consistent improvements across various tasks, including general VQA (e.g., 77.1% on MMB and 49.4% on MMStar), ultra-high-resolution image perception (e.g., 59.9% on HR-Bench). Notably, LLaVA-UHD v2 surpasses OCR-centric models like TextMonkey on DocVQA (72.9% vs. 66.5%) and outperforms those with multiple experts (such as MG-LLaVA), achieving superior performance on general tasks like SEED (73.6%). These results underscore the value of rich semantics derived from multi-level multi-modal granularity, enhancing both the understanding and perception abilities of MLLMs. (2) LLaVA-UHD v2 indicates efficiency on data utilization and computation. Compared to LLaVA-Next and Mini-Gemini, both operating at a 672$\times$672 resolution, LLaVA-UHD v2 supports 1.5 times the resolution ($i.e$., 672$\times$1008) and achieves superior performance with less than 40% of the computational cost. Furthermore, in contrast to Honey-bee and VILA, which utilize 52.5M and 51M data samples respectively, LLaVA-UHD v2 attains comparable or superior performance using only $\sim$2.8of the data, demonstrating the data efficiency of our model. As for the training duration, under the same model configuration and data volume, LLaVA-UHD v2 requires $\sim$27 hours to train on 8$\times$×A100 GPUs, while LLaVA-Next needs $\sim$42 hours, which is well-suited for low-cost exploratory research in the academic community.
4.4 Analytical Study
We conduct analytical experiments on the proposed modules to verify the effect of LLaVA-UHD v2. Without special instructions, we use Vicuna-7B as the base LLM.
Main module ablation. In Table 2, by replacing the low-resolution CLIP-ViT features with highest-level ones of the inverse semantic pyramid (ISP) constructed by VDIM, 2.2% average improvement can be seen, especially, on tasks depending on visual details like OCR-Bench (+9.5%) and RefCOCOs (+4.0%). Employing the Hiwin attention to integrate the ISP further increases 1.3% in average accuracy, especially 5.0% on ultra high-resolution perception (HR-Bench), demonstrating rich visual granularity could facilitate precise language generation. With spatially consistent token organization, visual-spatial understanding observes further improvement, such as 1.3% and 0.6% on RealWorldQA and RefCOCOs, respectively.
Table 3: Comparison of different methods for semantic pyramid construction. “ConvNext" means we replace the CILP-ViT with CLIP-ConvNext [73] as visual encoder and directly use the feature maps from multiple stages as the final hierarchical feature pyramid.